# The 18 May 2024 superbolide over the Iberian Peninsula: USG space sensors and ground-based independent observations

### v2

@@ Abstract
On 18 May 2024, a superbolide traversed the western part of the Iberian Peninsula, culminating its flight over the Atlantic Ocean and generating significant media attention. This event was caused by a weak carbonaceous meteoroid of 89.8$\pm$0.4 cm, with a density of 1660 kg m-3, entering the atmosphere at 41.6 km s-1 with an angle of 10.9. The luminous phase started at 138 km and ended at an altitude of 54 km. The meteoroid’s heliocentric orbit was characterized by an inclination of 15.4, a high eccentricity of 0.965, a semi-major axis of 3 au, and a notably short perihelion distance of 0.11 au. The superbolide was recorded by multiple ground-based stations of the Spanish Meteor Network (SPMN), the European Space Agency (ESA), and the U.S. Government (USG) space sensors. Our analysis shows a relatively good agreement with the radiant and velocity data reported by the Center for Near-Earth Object Studies (CNEOS), with a deviation of 1 and 1.2 km s-1, respectively. Due to the absence of observable deceleration, we successfully reconciled satellite radiometric data with a purely dynamic atmospheric flight model, constraining the meteoroid’s mass and coherently fitting its velocity profile. The inferred bulk density and aerodynamic strength of the meteoroid from the flight model are compatible with the properties recently measured for sample-returned materials from asteroid Bennu. The physical properties and the orbital results suggest that this meteoroid originated from a recent disruption of a comet, indicating the existence of hazardous meter-sized projectiles arriving on Earth from objects formed in the distant regions of the Solar System.
@@
Superbolide – Fireball Network – Space Sensors – Atmospheric Flight – Orbit
pubyear: 2024pagerange: The 18 May 2024 superbolide over the Iberian Peninsula: USG space sensors and ground-based independent observationsThe 18 May 2024 superbolide over the Iberian Peninsula: USG space sensors and ground-based independent observations
@@ 1 Introduction
On 18 May 2024, an exceptionally luminous fireball was observed over Spain and Portugal. This event was captured on video by casual observers and quickly disseminated through the media. Additionally, it was recorded by various wide-field and multi-camera stations that continuously monitor the sky over the Iberian Peninsula. The U.S. Government (USG) space sensors also detected the event, as reported on the Center for Near-Earth Object Studies (CNEOS) fireball website [^1] https://cneos.jpl.nasa.gov/. Following confirmation of its detection from space [^2] It was also detected by ESA’s Meteosat 3Gen (MTG) satellite via its Lightning Imager instrument: https://www.esa.int/Applications/Observing_the_Earth/Meteorological_missions/meteosat_third_generation/Fireball_witnessed_by_weather_satellite, the event can be formally classified as a superbolide (Ceplecha et al., 1999). These exceptionally bright fireballs are produced by the hypersonic atmospheric entry of meter-sized natural projectiles (Ceplecha et al., 1998; Silber et al., 2018). The study of superbolides provides valuable insights into the physical properties, dynamics, and impact hazard issues associated with the near-Earth asteroid population (Koschny & Borovicka, 2017; Trigo-Rodríguez, 2022).
@@
The CNEOS catalog, managed by NASA’s Jet Propulsion Laboratory, archives data on fireball events detected via various USG satellite sensors (Tagliaferri et al., 1994). Superbolides are relatively rare events that can occur over remote areas. USG sensors provide near-global coverage of large meteoroids and small asteroids impacting the Earth’s atmosphere, unlike the ground-based meteor network which only monitors a relatively small, fixed atmospheric volume. Therefore, the CNEOS catalog is of scientific interest as it extends the projectile flux estimations by utilizing the entire planet as a detector, capturing events that are typically singular occurrences for other techniques (Brown et al., 2002).
As of May 2024, the catalog contains 979 fireball events, with velocity vector and altitude data available for 310 cases. For these events, detailed information is provided, including geographic coordinates, terminal disruption altitude, vector velocity components in an Earth-Centered, Earth-Fixed (ECEF) reference frame, and estimated radiated and impact energies. These energies, linked by an empirical formula, are among the most reliable parameters provided by CNEOS (Brown et al., 2002).
@@
However, specific information about these sensors remains undisclosed, as they are classified data. In any case, we previously prepared our 3D-FireTOC software to compute future detections based on CNEOS-released data (Peña-Asensio et al., 2021b) and performed an analysis of the CNEOS catalog (Peña-Asensio et al., 2022). The accuracy of the CNEOS data has been assessed through comparisons with ground-based observations of fireballs (Devillepoix et al., 2019; Peña-Asensio et al., 2022; Brown & Borovička, 2023; Peña-Asensio et al., 2024). To date, only 17 fireballs in the CNEOS catalog have been benchmarked with independent counterparts. Here we report and compare a new event detected by the USG space sensors and multiple ground-based fireball networks.
@@ 2.1 Ground-based observation
On the night of Saturday, May 18, 2024, a remarkable superbolide was observed over the Iberian Peninsula at 22:46:49 UTC, specifically flying over Extremadura and northern Portugal, before its trajectory concluded over the Atlantic Ocean. It was recorded by the Spanish Fireball and Meteorite Network (SPMN) (Trigo-Rodríguez et al., 2004), as well as by AMS82, one of the European Space Agency’s (ESA) Planetary Defense Office meteor stations of the AllSky7 Network333www.allsky7.com, located in Casas de Millán, Cáceres, Spain. Designated as SPMN180524F by the SPMN network, this event was notable for its intense brightness which momentarily turned night into day, causing a massive media impact. The superbolide exhibited an atmospherically originated bluish-green glow and left a persistent luminous trail in the sky. Table 1 lists the stations used in this study, while Figure 1 shows the max-combined video frames as observed from each station.
Many casual videos immediately surfaced in the media, underscoring the importance of promptly explaining these unusual phenomena to the public. Some of the videos are particularly extraordinary, as the bolide was captured from a relatively close distance, revealing significant variations in brightness along its extensive luminous trajectory. These fluctuations in luminosity are typically associated with the continuous fragmentation of the meteoroid and the release of dust and fragments ablated by the heat generated in the frontal shock wave, a characteristic behavior of impacting crumbly meteoroids (Revelle, 2002).
Figure 1: The four videos used in this work to study the superbolide SPMN180524F. The images are max-combined from the video frames. Some saturated frames have been removed for illustration purposes. From left to right: Navianos de Valverde, Estepa, Sanlúcar de Barrameda, and Casas de Millán.
@@
Navianos de Valverde 5 48 48.4"""" W 41 57 11.6"""" N 711
Estepa 4 52 35.6"""" W 37 17 29"""" N 537
Sanlúcar de Barrameda 6 20 31"""" W 36 46 30.3"""" N 6
Casas de Millán 6 19 57.8"""" W 39 51 4.8"""" N 739
We use our 3D-FireTOC software to analyze the event. This tool integrates computer vision and machine learning for detecting and tracking fireballs, robust camera calibration to correct distortions, the plane intersection method for triangulation, and heliocentric orbit determination (Peña-Asensio et al., 2021b, a, 2023b, 2023a). Using a model of atmospheric mass density and the measured velocity, the dynamic strength where the meteoroid disruption occurs is computed as $S=\rho_{atm}v^{2}$ (Bronshten, 1983). From that, thanks to the empirical strength-density relation established by Collins et al. (2005), we calculate the impactor density as:
$\rho_{m}=\left(\frac{\log_{10}S-2.107}{0.0624}\right)^{2}.$ (1)
At the onset of the luminous phase and down to altitudes near 60 km, the atmosphere is sufficiently thin that deceleration is barely observable. Consequently, all measured points during this phase equally represent the initial velocity, allowing for a robust estimate.
@@ 2.2 Space-based observation
From the perspective of the CNEOS catalog, the SPMN180524F superbolide, with a total impact energy of 0.13 kt, is not a common event: it holds the highest height in the database, ranks as the fifth fastest with respect to the Earth, and is the fourth with the lowest density (see Fig. 2). The density estimation process is the same as described in the previous section.
Figure 2: CNEOS fireballs with sufficient data (velocity and height at the peak of radiated energy) to estimate the projectile density.
As the superbolide did not penetrate deeply into the atmosphere and was observed from long distances, the ground-based measurements are insufficient for fitting a dynamic deceleration model. Therefore, we use as a proxy the total radiated energy ($E$) measured by the USG sensors, which has proven reliable. By assuming that the energy recorded is equal to the kinetic energy, we estimate the initial mass of the projectile as:
@@
$\alpha=\frac{c_{d}\rho_{0}h_{0}A}{2m_{0}^{1/3}\rho_{m}^{2/3}\sin\gamma},$ (3)
where $c_{d}=0.7$ is the drag coefficient, $A=1.21$ is the spherical shape factor, $\rho_{0}=1.29\,\,$kg m-3 is the atmospheric density at the sea level, $h_{0}=7.16\,\,$km is the height of the homogeneous atmosphere, and $\gamma$ is the slope of the trajectory from the horizon.
By assuming the terminal mass is equal to zero, $\beta$ can be estimated from the terminal height $h_{e}$ and $\alpha$ (Moreno-Ibáñez et al., 2015):
@@
3 Results and comparison
Figure 3 displays the 3D reconstruction of the atmospheric flight, and Figure 4 depicts its characterization. The heliocentric orbits derived from the ground-based observations and the space sensors are illustrated in Figure 5. Table 2 shows all parameters derived from the ground- and space-based observations.
Figure 3: Atmospheric flight of the superbolide SPMN180524F. For ground-based observations, the peak brightness occurs at the red point. In pink is shown the point reported by the USG space sensors.
Figure 4: Characterization of the atmospheric flight of the superbolide SPMN180524F. It includes the measured velocity points, the best fit for the velocity with 3-$\sigma$ uncertainty, the dynamic strength, and the uncalibrated photometric counts from Estepa.
Figure 5: Osculating heliocentric orbit of the superbolide SPMN180524F determined from ground and space observations.
Table 2: Atmospheric flight, physical parameters, and osculating heliocentric orbital elements of the SPMN180524F superbolide.
@@ 3 Results and comparison
Initial time [UTC] $t_{0}$ 2024-05-18 22:46:42
Initial longitude [] $\lambda_{0}$ 5.5197$\pm$0.0010 W
Initial latitude [] $\varphi_{0}$ 38.5777$\pm$0.0011 N
Initial velocity [km s-1] $v_{0}$ 41.64$\pm$0.19
Initial height [km] $h_{0}$ 137.88$\pm$0.05
Final time [UTC] $t_{e}$ 2024-05-18 22:46:53
Final longitude [] $\lambda_{e}$ 9.885$\pm$0.005 W
Final latitude [] $\varphi_{e}$ 42.112$\pm$0.005 N
Final velocity [km s-1] $v_{e}$ 40.03$\pm$0.17
Final height [km] $h_{e}$ 53.78$\pm$0.07
Energy peak time [UTC] $t_{p}$ 2024-05-18 22:46:49 2024-05-18 22:46:50 -1 s
Energy peak longitude [] $\lambda_{p}$ 8.2932$\pm$0.0017 W 8.8 W -0.5068
Energy peak latitude [] $\varphi_{p}$ 40.8851$\pm$0.0015 N 41.0 N -0.1149
Energy peak velocity [km s-1] $v_{p}$ 41.63$\pm$0.19 40.4 1.23
Energy peak height [km] $h_{p}$ 77.91$\pm$0.04 74.3 3.61
Length [km] $\Delta l$ 554.2$\pm$0.4
Slope [] $\gamma$ 10.93$\pm$0.02 6.5 4.43
Azimuth [] $A$ 317.75$\pm$0.04 315.8 1.95
Dynamic strength [kPa] $S$ 44.4$\pm$0.4 72.41 -28.01
Meteoroid density [kg m-3] $\rho_{m}$ 1658$\pm$6 1946 -288
Meteoroid mass [kg] $m_{0}$ 628$\pm$6 666 -38
Initial diameter [cm] $D$ 89.8$\pm$0.4 87 2.8
Geo. velocity [km s-1] $v_{R}$ 39.87$\pm$0.19 38.6 1.27
Geo. radiant R.A. [] $\alpha_{R}$ 261.56$\pm$0.04 262.4 -0.84
Geo. radiant Dec. [] $\delta_{R}$28.13$\pm$0.04 -29.7 1.57
Semi-major axis [au] $a$ 3.01$\pm$0.11 2.3 0.71
Eccentricity $e$ 0.965$\pm$0.002 0.95 0.015
Inclination [] $i$ 15.39$\pm$0.12 18.2 -2.81
Argument of periapsis [] $\omega$ 145.52$\pm$0.12 145.4 0.12
Long. of the asc. node [] $\Omega$ 238.1257$\pm$0.0001 238.1321 -0.0064
Perihelion distance [au] $q$ 0.106$\pm$0.001 0.11 -0.004
True anomaly [] $f$ 214.49$\pm$0.12 214.6 -0.11
Period [year] $P$ 5.2$\pm$0.3 3.5 1.7
Tisserand’s parameter $Tj$ 2.1$\pm$0.1 2.7 -0.6
Ballistic coefficient $\alpha$ 17.22$\pm$0.05
Mass-loss parameter $\beta$ 53.1$\pm$0.4
The search for possible associations with meteoroid streams or parent bodies for superbolide SPMN180524F yielded negative results using both traditional D-criteria and Machine Learning distance metrics (Peña-Asensio & Sánchez-Lozano, 2024). If the SPMN180524F superbolide resulted from a cometary disruption, additional fragments could potentially reach the Earth. Nevertheless, these are difficult to detect in advance due to the typical low albedo of carbonaceous chondrites (Trigo-Rodríguez et al., 2014; Tanbakouei et al., 2020). Given the estimated meteoroid size of approximately 0.9 m and a density of 1660 kg m-3, we hypothesize its origin might be the disruption of a comet. Remarkably, a recent study of the samples returned by the OSIRIS-REx NASA mission revealed similar bulk density values (Lauretta et al., 2024), which could suggest a Bennu-like parent body.

### v4

# The 18 May 2024 Iberian superbolide from a sunskirting orbit: USG space sensors and ground-based independent observations

@@ Abstract
On 18 May 2024, a superbolide traversed the western part of the Iberian Peninsula, culminating its flight over the Atlantic Ocean and generating significant media attention. This event was caused by a weak carbonaceous meteoroid of 1 m, entering the atmosphere at 40.4 km s-1 with an average slope of 8.5. The luminous phase started at 133 km and ended at an altitude of 54 km. The meteoroid’s heliocentric orbit had an inclination of 16.4, a high eccentricity of 0.952, a semi-major axis of 2.4 au, and a short perihelion distance of 0.12 au. The superbolide was recorded by multiple ground-based stations of the Spanish Fireball and Meteorite Network (SPMN) and the European Space Agency (ESA), as well as by the U.S. Government (USG) sensors from space. Due to the absence of observable deceleration, we successfully reconciled satellite radiometric data with a purely dynamic atmospheric flight model, constraining the meteoroid’s mass and coherently fitting its velocity profile. Our analysis shows a good agreement with the radiant and velocity data reported by the Center for Near-Earth Object Studies (CNEOS), with a deviation of 0.56 and 0.1 km s-1, respectively. The presence of detached fragments in the lower part of the luminous trajectory suggests that the meteoroid was a polymict carbonaceous chondrite, containing higher-strength macroscopic particles in its interior due to collisional gardening, or a thermally processed C-type asteroid. The orbital elements indicate that the most likely source is the Jupiter-Family Comet (JFC) region, aligning with the SOHO comet family, as its sunskirting orbit is decoupled from Jupiter. This event provides important information to characterize the disruption mechanism of near-Sun objects.
@@
meteorites, meteors, meteoroids – comets: general – minor planets, asteroids: general
pubyear: 2024pagerange: The 18 May 2024 Iberian superbolide from a sunskirting orbit: USG space sensors and ground-based independent observationsThe 18 May 2024 Iberian superbolide from a sunskirting orbit: USG space sensors and ground-based independent observations
@@ 1 Introduction
On 18 May 2024, an exceptionally luminous fireball was observed over Spain and Portugal. This event was captured on video by casual observers and quickly disseminated through the media. Additionally, it was recorded by various wide-field and multi-camera stations that continuously monitor the sky over the Iberian Peninsula. The U.S. Government (USG) sensors also detected the event from space, as reported on the Center for Near-Earth Object Studies (CNEOS) fireball website [^1] https://cneos.jpl.nasa.gov/. Following confirmation of its detection from space [^2] It was also detected by ESA’s Meteosat 3Gen (MTG) satellite via its Lightning Imager instrument: https://www.esa.int/Applications/Observing_the_Earth/Meteorological_missions/meteosat_third_generation/Fireball_witnessed_by_weather_satellite, the event can be formally classified as a superbolide (Ceplecha et al., 1999). These exceptionally bright fireballs are produced by the hypersonic atmospheric entry of meter-sized natural projectiles (Ceplecha et al., 1998; Silber et al., 2018). The study of superbolides provides valuable insights into the physical properties, dynamics, and impact hazard issues associated with the near-Earth object population (Koschny & Borovicka, 2017; Trigo-Rodríguez, 2022).
@@
The CNEOS database, managed by NASA’s Jet Propulsion Laboratory, archives data on fireball events detected via various USG satellite sensors (Tagliaferri et al., 1994). Superbolides are relatively rare events that can occur over remote areas. USG sensors provide near-global coverage of large meteoroids and small asteroids impacting the Earth’s atmosphere, unlike the ground-based meteor network which only monitors a relatively small, fixed atmospheric volume. Therefore, the CNEOS database is of scientific interest as it extends the projectile flux estimations by utilizing the entire planet as a detector, capturing events that are typically singular occurrences for other techniques (Brown et al., 2002). As of May 2024, the database contains 979 fireball events, with velocity vector and position data available for 310 cases. For these events, detailed information is provided, including geographic coordinates and altitude of the radiated energy peak, vector velocity components in an Earth-Centered, Earth-Fixed (ECEF) reference frame, and radiated and impact energies. These energies, linked by an empirical formula, are among the most robust parameters provided by CNEOS (Brown et al., 2002).
@@
However, specific information about these sensors remains undisclosed as they are classified data. In any case, we previously prepared our 3D-FireTOC software to compute future detections based on CNEOS-released data (Peña-Asensio et al., 2021b) and performed an analysis of the CNEOS database (Peña-Asensio et al., 2022). The accuracy of the CNEOS data has been assessed through comparisons with ground-based observations of fireballs (Devillepoix et al., 2019; Peña-Asensio et al., 2022; Brown & Borovička, 2023; Peña-Asensio et al., 2024). To date, only 17 fireballs in the CNEOS database have been benchmarked with counterparts. Here we report and compare a new event detected by the USG sensors and multiple ground-based networks.
@@ 2.1 Ground-based observation
On the night of Saturday, May 18, 2024, a remarkable superbolide was observed over the Iberian Peninsula at 22:46:48 UTC, specifically crossing Extremadura and northern Portugal before its flight concluded over the Atlantic Ocean. It was recorded by the Spanish Fireball and Meteorite Network (SPMN, Trigo-Rodríguez et al., 2004), as well as by AMS82, one of the European Space Agency’s (ESA) Planetary Defense Office meteor stations of the AllSky7 Network333www.allsky7.com, located in Casas de Millán, Cáceres, Spain. Designated as SPMN180524F by the SPMN network and popularized as the 2024 Iberian Superbolide, this event was notable for its intense brightness which momentarily turned night into day, causing a massive media impact. The superbolide exhibited an atmospherically originated bluish-green glow, leaving a persistent luminous trail in the sky. Table 1 lists the stations used in this study, while Figure 1 shows the max-combined video frames observed from each station.
Many casual videos immediately surfaced in the media, underscoring the importance of promptly explaining these unusual phenomena to the public. Some of the videos are particularly extraordinary, as the bolide was captured from close distances, revealing significant variations in brightness along its extensive luminous trajectory. These fluctuations in luminosity are typically associated with the continuous fragmentation of the meteoroid and the release of dust and fragments ablated by the heat generated in the frontal shock wave, a characteristic behavior of impacting crumbly meteoroids (Revelle, 2002). In some videos of the final part of the luminous phase, the bolide head is followed by distinct ablating fragments (see Fig. 1).
Figure 1: The four videos used in this work to study the SPMN180524F superbolide. The images are max-combined from the video frames. Some saturated frames have been removed for illustration purposes. From left to right: Navianos de Valverde, Estepa, Sanlúcar de Barrameda, and Casas de Millán. At the center is a frame from a video shared on social media that captures the final part of the luminous phase.
@@
Navianos de Valverde 5 48 48.4′′ W 41 57 11.6′′ N 711
Estepa 4 52 35.6′′ W 37 17 29.0′′ N 537
Sanlúcar de Barrameda 6 20 31.0′′ W 36 46 30.3′′ N 6
Casas de Millán 6 19 43.0′′ W 39 48 57.3′′ N 456
We calibrate the cameras using the SkyFit program provided in the RMS library, which incorporates a 7th-order polynomial distortion model and accounts for atmospheric refraction (Vida et al., 2021). We use our 3D-FireTOC software to analyze the event. This tool integrates the plane intersection method for trajectory triangulation and an N-body numerical integrator for heliocentric orbit determination (Peña-Asensio et al., 2021b, a, 2023b, 2023a). Using a model of atmospheric mass density and the measured velocity, the dynamic strength where the meteoroid disruption occurs is computed as $S=\rho_{atm}v^{2}$ (Bronshten, 1983). A traditional method for classifying fireballs is the $P_{E}$ criterion established by Ceplecha & McCrosky (1976), which serves to evaluate physical properties of meteoroids. Inspired by this criterion, Borovička et al. (2022b) introduced a new parameter for assessing the impactor strength, based on the maximum dynamic pressure for large meteoroids. This parameter, termed the pressure resistance factor or pressure factor, is defined as follows:
$Pf=100S_{max}\sin{\overline{\gamma}}^{-1}m_{0}^{-1/3}v_{0}^{-3/2},$ (1)
where $S_{max}$ represents the maximal dynamic pressure in MPa, $\overline{\gamma}$ is the average slope of the trajectory from the local horizon, $m_{0}$ denotes the initial photometric mass in kg, and $v_{0}$ is the entry velocity in km s-1.
At the start of the luminous phase and down to altitudes near 60 km, the atmosphere is not dense enough to cause observable deceleration. Consequently, all measured points during this phase equally represent the initial velocity, allowing for a robust estimate despite the point-by-point measurement dispersion.
@@ 2.2 Space-based observation
From the perspective of the CNEOS database, the SPMN180524F superbolide, with a total impact energy of 0.13 kt, is not a common event: it holds the highest altitude in the database, ranks as the fifth fastest with respect to the Earth, and is the fourth with the lowest dynamic strength (see Fig. 2). We estimated the CNEOS fireball density in the same way as described in the previous section.
Figure 2: CNEOS fireballs with sufficient data (velocity and height at the peak of radiated energy) to estimate the projectile dynamic strength.
As the superbolide did not penetrate deeply into the atmosphere and was observed from long distances, the ground-based measurements are insufficient for fitting a dynamic deceleration model. Therefore, we use as a proxy the total impact energy ($E$) measured by the USG sensors, which has proven reliable. By assuming that the energy recorded is equal to the kinetic energy, we estimate the initial mass of the projectile as:
@@
$\alpha=\frac{c_{d}\rho_{sl}h_{h}A}{2m_{0}^{1/3}\rho_{m}^{2/3}\sin\overline{\gamma}},$ (3)
where $c_{d}=0.7$ is the drag coefficient, $A=1.21$ is the spherical shape factor, $\rho_{sl}=1.29\,\,$kg m-3 is the atmospheric density at the sea level, $\rho_{m}$ is the meteoroid density, and $h_{h}=7.16\,\,$km is the height of the homogeneous atmosphere. By assuming the final mass is equal to zero, $\beta$ can be estimated from the final height $h_{e}$ and $\alpha$ (Moreno-Ibáñez et al., 2015):
@@
To search for possible associations with specific parent bodies or meteoroid streams, we use both traditional D-criteria and Machine Learning distance metrics (Peña-Asensio & Sánchez-Lozano, 2024). We employ the NEOMOD model to explore the possible source of the heliocentric orbit (Nesvorný et al., 2023, 2024a, 2024b). NEOMOD integrates asteroid orbits from main belt sources and comets, calibrates results with observations from the Catalina Sky Survey, and includes visible albedo information from the Wide-Field Infrared Survey Explorer (WISE).
3 Results and discussion
3.1 Atmospheric flight, physical properties, and orbit
Figure 3 displays the 3D reconstruction of the atmospheric flight, and Figure 4 depicts its velocity as a function of height, along with the dynamic strength and the uncalibrated photometric count. The heliocentric orbits derived from ground- and space-based observations are illustrated in Figure 5. Table 2 shows all derived parameters and their comparison where possible.
Figure 3: Atmospheric flight of the SPMN180524F superbolide. For ground-based observations, the peak brightness occurs at the red point. In pink is shown the point reported by the USG sensors.
Figure 4: Characterization of the atmospheric flight of the SPMN180524F superbolide. It includes the measured velocity points, the best fit for the velocity with 3-$\sigma$ uncertainty, the dynamic strength, and the uncalibrated photometric counts from Estepa. The section without measured points corresponds to oversaturated frames.
Figure 5: Osculating heliocentric orbit of the SPMN180524F superbolide determined from ground- and space-based observations.
Table 2: Atmospheric flight, physical parameters, and osculating heliocentric orbital elements of the SPMN180524F superbolide. aAssuming the radiated energy reported by CNEOS is perfectly accurate. bOur value represents the average, whereas CNEOS’s value corresponds to the energy peak.
@@ 3.1 Atmospheric flight, physical properties, and orbit
Initial time [UTC] $t_{0}$ 2024-05-18 22:46:41
Initial longitude [] $\lambda_{0}$ 5.4608$\pm$0.0003 W
Initial latitude [] $\varphi_{0}$ 38.4738$\pm$0.0011 N
Initial velocity [km s-1] $v_{0}$ 40.4$\pm$0.4
Initial height [km] $h_{0}$ 132.96$\pm$0.16
Energy peak time [UTC] $t_{p}$ 2024-05-18 22:46:48 2024-05-18 22:46:50 -2 s
Energy peak longitude [] $\lambda_{p}$ 8.28$\pm$0.11 W 8.8 W -0.52
Energy peak latitude [] $\varphi_{p}$ 40.85$\pm$0.09 N 41.0 N -0.15
Energy peak velocity [km s-1] $v_{p}$ 40.3$\pm$0.4 40.4 -0.1
Energy peak height [km] $h_{p}$ 74.5$\pm$1.8 74.3 0.2
Final time [UTC] $t_{e}$ 2024-05-18 22:46:55
Final longitude [] $\lambda_{e}$ 9.724$\pm$0.006 W
Final latitude [] $\varphi_{e}$ 41.989$\pm$0.007 N
Final velocity [km s-1] $v_{e}$ 38.3$\pm$0.4
Final height [km] $h_{e}$ 53.59$\pm$0.24
Length [km] $\Delta l$ 546.2$\pm$0.9
[…]
Dynamic strength [kPa] $S$ 73$\pm$19 72.4 0.6
[…]
Meteoroid density [kg m-3] $\rho_{m}$ 1500$\pm$300
Initial diameter [m] $D$ 1.0$\pm$0.1
Geo. velocity [km s-1] $v_{R}$ 38.5$\pm$0.4 38.6 -0.1
Geo. radiant R.A. [] $\alpha_{R}$ 261.79$\pm$0.03 262.4 -0.61
Geo. radiant Dec. [] $\delta_{R}$29.11$\pm$0.07 -29.7 0.59
Semi-major axis [au] $a$ 2.43$\pm$0.15 2.3 0.13
Eccentricity $e$ 0.952$\pm$0.004 0.95 0.002
Inclination [] $i$ 16.36$\pm$0.31 18.2 -1.84
Argument of perihelion [] $\omega$ 144.64$\pm$0.14 145.4 -0.76
Long. of the asc. node [] $\Omega$ 238.1262$\pm$0.0001 238.1321 -0.0059
Perihelion distance [au] $q$ 0.116$\pm$0.003 0.11 0.006
True anomaly [] $f$ 215.36$\pm$0.14 214.6 0.76
Period [year] $P$ 3.79$\pm$0.34 3.5 0.29
Tisserand’s parameter $T_{J}$ 2.55$\pm$0.13 2.66 -0.11
Ballistic coefficient $\alpha$ 23.6$\pm$3.1
Mass-loss parameter $\beta$ 38$\pm$5
For the SPMN180524F superbolide, we computed a $P_{E}$ value of -5.15, which is slightly above the threshold of -5.25 that differentiates between carbonaceous and regular cometary material types. To contextualize this event, we compared it to the catalog of 824 fireballs recorded and analyzed by the European Fireball Network (EN) (Borovička et al., 2022a, b). We calculated a $Pf$ value of 0.31 for SPMN180524F, placing it in class $Pf$-II. Meteoroids with $Pf$ values less than 0.27 are considered cometary, while those with $Pf$ greater than 0.85 are considered asteroidal. Therefore, we interpret SPMN180524F as a weak carbonaceous body and draw its bulk density from a uniform distribution between 1000 and 2000 kg m-3 for subsequent calculations. Figure 6 shows the comparison with the EN catalog. The SPMN180524F superbolide appears as an outlier in the final versus initial height panel, possibly due to its grazing atmospheric slope and size, as the largest mass in the catalog is 110 kg. This event is not particularly unusual when considering the combination of the Tisserand parameter and the pressure factor, as similar carbonaceous impactors are found in Jupiter-Family Comet (JFC) orbits. SPMN180524F stands out due to its perihelion distance (0.12 au), which is within the shortest in the catalog. In the eccentricity versus semi-major axis panel, the SPMN180524F event appears grouped with other $Pf$-II and $Pf$-I events, lying well within the JFC region and approaching the so-called excited short-period orbits (Borovička et al., 2022b). SPMN180524F is the second most pressure-resistant $Pf$-II event. Remarkably, the samples returned by the OSIRIS-REx NASA mission revealed similar bulk density values for the carbonaceous asteroid Bennu (Lauretta et al., 2024). We note that carbonaceous chondrite (CC) projectiles are difficult to detect before the atmospheric impact due to their typical low albedo (Trigo-Rodríguez et al., 2014; Tanbakouei et al., 2020)
Figure 6: Comparison of the EN catalog and the SPMN180524F superbolide. Top left: final and initial heights, color-coded by initial velocity. Top right: height at maximum pressure and maximum pressure, classified by the pressure factor. Bottom left: pressure factor and Tisserand parameter with respect to Jupiter, color-coded by perihelion distance. Bottom right: eccentricity and semi-major axis distribution, classified by the pressure factor. Asteroids and comets with perihelion distances lower than 1 au are plotted. The inner white edge denotes fireballs that originate with 100% confidence from a JFC source (see Table 3). Comets with an orange edge represent the SOHO family, and those with a pink edge are from the Machholz complex.
Searching for associations with meteoroid streams or parent bodies yielded negative results. Nevertheless, $\epsilon$-Scorpiids meteor shower exhibits the highest similarity with $D_{D}$ = 0.08 (Drummond, 1981). By estimating the source contributions to the major axis, eccentricity, inclination, and absolute magnitude space, NEOMOD determined that the orbit of the SPMN180524F impactor originated from a JFC with a 100% confidence. On the other hand, while the semi-major axis of $\sim$2.5 au might suggest an origin from the 3:1 resonance, this hypothesis is not supported by NEOMOD and should be verified numerically. We emphasize that the dynamic transit of an asteroid from the outer main asteroid belt cannot be ruled out due to the chaotic nature of these processes (Valsecchi et al., 1995). We have also run the NEOMOD model on the EN catalog and listed in Table 3 the events with 100% confidence of having a JFC source. Among these, SPMN180524F was the only meter-sized impactor, had the highest initial altitude by a significant margin and the shallowest atmospheric entry angle, and experienced the greatest dynamic strength. Note that NEOMOD can extrapolate results for two-thirds of the catalog, and only 7.6% of events with $T_{J}$ between 2 and 3 have over a 95% probability of originating from a JFC source.
Table 3: List of EN fireballs with orbits that NEOMOD identifies as 100% originating from JFC sources. The superbolide SPMN180524F is included for comparison. The events are sorted by photometric mass.
Event $m_{0}$ $h_{0}$ $h_{e}$ $\overline{\gamma}$ $v_{0}$ $S_{max}$ $Pf$ class $T_{J}$ $q$ $e$ $i$
[kg] [km] [km] [] [km s-1] [MPa] [au] []
SPMN180524F 670 132.96 53.59 40.4 1.015 $Pf$-II 2.55 0.116 0.952 16.36
EN020617_230339 0.57 97.37 52.80 12.56 36.993 0.434 $Pf$-I 2.511 0.2017 0.9234 11.52
EN020818_012951 0.45 101.62 53.04 39.87 41.273 0.819 $Pf$-II 2.043 0.1052 0.9661 22.67
EN040818_014525 0.035 100.45 66.95 25.24 41.862 0.177 $Pf$-II 1.976 0.0937 0.9707 22.96
EN180517_013552 0.015 92.38 60.52 25.68 39.743 0.302 $Pf$-I 2.501 0.1333 0.9467 19.69
EN040717_223844 0.011 95.06 51.30 38.92 41.700 0.755 $Pf$-I 1.890 0.2163 0.9372 48.15
EN300717_021036 0.0072 98.31 69.04 23.07 42.998 0.189 $Pf$-I 1.890 0.0757 0.9769 27.50
EN160817_004534 0.005 88.68 67.88 32.11 42.119 0.181 $Pf$-II 2.220 0.0757 0.9722 22.47
EN080818_005515 0.0018 91.93 70.30 27.49 41.386 0.133 $Pf$-I 2.068 0.1067 0.9650 25.00
EN180717_000606 0.001 101.31 79.05 16.73 41.098 0.037 $Pf$-II 2.440 0.1158 0.9534 33.60
EN161118_233946 0.00083 99.36 66.96 66.2 41.093 0.142 $Pf$-II 2.365 0.0944 0.9638 8.49
EN061218_013732 0.00052 92.64 63.04 67.39 42.235 0.203 $Pf$-I 2.438 0.0782 0.9682 20.60
EN091018_014119 0.00037 100.69 63.52 51.10 37.664 0.140 $Pf$-I 2.247 0.1740 0.9419 4.61
EN081118_011907 0.00015 94.78 71.09 55.34 42.043 0.073 $Pf$-II 2.409 0.0844 0.9661 26.38
EN060818_212305 0.00013 97.26 87.75 13.11 41.819 0.008 $Pf$-III 2.083 0.1007 0.9665 25.44